(i) Compute the value of 𝛾 for a particle traveling at half the speed of light. Give your answer


to three significant figures.


(ii) Determine the energy required to give an electron a speed of 0.90 that of light, starting from


rest.

1
Expert's answer
2022-02-11T14:17:26-0500

(i) We find the value for gamma with

"γ=\frac{1}{\sqrt{1−(v/c)²}}
\\ \gamma =\frac{1}{\sqrt{1−(0.500)²}}=\frac{1}{\sqrt{0.750}}
\\ \therefore \gamma =\frac{1}{0.866}=1.15"


(ii) The kinetic energy required is equal to


"KE = (\gamma m - m)c²=(\gamma - 1)mc²
\\ \therefore KE= mc² \Big(\frac{1}{\sqrt{1−(v/c)²}}-1\Big)
\\ \text{We substitute and find KE}
\\ KE = (\frac{1}{\sqrt{1−(0.90)²}}-1)(9.11×10^{-31}{kg})(3×10⁸\frac{m}{s})²
\\ \therefore KE = 1.06×10^{-13}J"


In conclusion, (i) 𝛾 = 1.15 when v/c = 0.50, while (ii) we need 1.06 × 10-13 J of kinetic energy to take an electron from the rest up to v/c = 0.90.